# Defects and Impurity Properties of VN precipitates in ARAFM Steels: Modelling using a Universal Machine Learning Potential and Experimental Validation

### v1

@@ Abstract
VN precipitates used to strengthen ARAFM steels for fusion applications, have been shown to undergo dissolution under high Fe ion irradiation doses of 100 dpa at dose rates of $10^{-3}$ dpa/s at 600 ${}^{o}C$. Here, point defects and solute substitutions have been studied using atom probe tomography (APT), universal machine learning interatomic potentials (uMLIPs), and density functional theory. Through a combination of transmission electron microscopy (TEM), APT, and atomic scale calculations, N-vacancies and substitutional Cr are found to be present in VN precipitates prior to irradiation. Ternary convex hulls were calculated for ten VNX (X=Cr, Fe, C, Si, Mn, W, Ta, B, S, P) systems with an uMLIP. These calculations predict Fe, P, Mn, and Si to be unstable solutes in the VN precipitate. Therefore, Fe, P, Mn and Si implantation via collision cascades is found to be a possible mechanism driving dissolution of VN.
@@
VN , RAFM , UMLIPs , APT
@@ 1 Introduction
One challenge for the commercialisation of fusion energy reactors is the development of reduced activation steels with high resistance to creep and irradiation damage that are easy to manufacture [1]. To this end, advanced reduced activation ferritic-martensitic (ARAFM) steel alloys with rocksalt-structured MX phase precipitates, such as vanadium nitride, have been proposed [2, 3, 4, 5]. However, Haley et al. recently observed that vanadium nitride precipitates underwent complete dissolution during a 100 dpa Fe+ ion irradiation [6]. ARAFM alloys are complex systems with multiple alloying elements in addition to precipitate phases. Thus, understanding the thermodynamic factors and role of irradiation damage in precipitate stability is both challenging and necessary.
Although alloy designers have often assumed MX precipitates to have idealized rocksalt stoichiometry and crystal structures, many rocksalt nitrides exhibit a strong preference for vacancy-driven non-stoichiometry [7, 8, 9, 10, 11, 12]. This preferred vacancy concentration may change if the MX phase is alloyed with other elements, either during precipitation or due to ballistic mixing. Additionally, during irradiation, the preferred vacancy concentration will be disturbed and other point defects will accumulate, which can further change the stability of the precipitate. A better understanding of the role of solutes and point defects on the stability and resilience of the VN precipitate under irradiation is needed.
We report a combined computational and experimental approach that sought to shed light on the structure of VN precipitates in ARAFM steel prior to irradiation and the types of irradiation damage that may cause dissolution. Atomic scale modelling, combined with TEM and APT results, pointed to a significant concentration of vacancies and solutes in the unirradiated precipitates. Ten V-N-X (X=Cr, Fe, C, Si, Mn, W, Ta, B, S, P) ternary systems of interest were identified as possible dissolution pathways primarily instigated by irradiation damage. Since, ternary systems represent a large undertaking for DFT calculations in terms of computational cost, we used a universal machine learning interatomic potential (UMLIP) to study these systems.
@@ 2 Methods
Samples of unirradiated ARAFM steel of the same composition and microstructure as studied under irradiation by Haley et al. were studied with TEM and APT. The lattice parameter of VN precipitates could be measured from the series of dark-field images collected by Haley et al. For each dark-field image, a corresponding diffraction pattern was captured, and the position of the objective aperture was recorded. These were compared against a simulated pattern of VN from which the lattice parameter was deduced. APT on the unirradiated ARAFM steel was collected via both voltage (20% pulse fraction, 200 kHz pulse frequency) and laser pulsing methods (40 pJ pulse energy, 200 kHz pulse frequency) at a temperature of 50 K. This data was analysed with APSuite 6.1. The precipitate compositions were measured using proxigrams created from 2.2% concentration isoconcentration surfaces with a bin size of 0.1 nm over three precipitates.
Density functional theory (DFT) calculations of point defects in VN were carried out using VASP 6 [13, 14, 15] using the Perdew, Burke, Ernzehof (PBE) exchange correlation functional [16] with k-point spacing and a planewave cut-off energy set to 0.2Å$\AA^{-1}$ and 520 eV, respectively. The projector augmented wave method (PAW) [17, 18] pseudopotentials with semicore p and s states were used for V, Fe, and Cr with Gaussian smearing of 0.1 eV for relaxation followed with a static run using the tetrahedron method with Blöchl corrections [19]. Atomic positions and unit cell parameters of all input structures were allowed to relax to a force convergence of 0.02 $/\AA$ per atom.
@@
The thermodynamic stability of the VN and VNX systems was studied by analysing the formation energy crystal structures with point defects and clusters of point defects. Thirty intrinsic point defects were calculated with DFT: V(N) vacancies, interstitials, and anti-sites. An additional 18 DFT calculations were performed for Cr and Fe substitutions on the V sublattice and C substitutions on the N sublattice. The ideal rocksalt unit cell has a single unique site for each defect type. To vary the defect concentration, a single defect was inserted in rocksalt supercells ranging in size from 8 atoms to 256 atoms. Configurations with multiple defects of different types were not considered.
Additional calculations were performed with a universal machine learning interatomic potential (UMLIP) called Orb-v1 [20] using the atom simulation environment (ASE)[21] implemented in python. For these calculations, each structure was relaxed using the Frechet cell filter, which allowed the box and atoms to relax. The fast inertial relaxation engine (FIRE) [22] was then applied with the convergence criteria of a maximum force component of 0.05 $/\AA$ on each atom or 1000 steps.
@@
Structures were calculated with the UMLIP as follows. First, an ideal conventional unit cell was initialized with ASE. Supercells were then enumerated using the integrated cluster expansion toolkit (ICET) [23]. The structure enumeration was broken into different series for each defect and substitution type. All possible unique structures containing the defect type were enumerated using ICET to a maximum of 14 atoms for the rocksalt structure. The Orb-v1 potential was used to calculate formation enthalpies for VN and 10 VNX systems, where X denotes the ten solute elements (Cr, Fe, C, Si, Mn, W, Ta, B, S, P). The defect types consisted of both V-N anti-sites, all three species interstitials, both V-N vacancies, and both V-N substitutions with the third species. While this was not an exhaustive search of all phases that could be included in each ternary system, it does include the rocksalt system and each element dissolved in BCC Fe, i.e. the two phases observed in experimental analysis.
@@
While formation energies are typically calculated relative to the elements standard crystal structures, we have chosen to use the energy of elements as solutes in body-centered cubic (BCC) iron. This was chosen because we are most interested in the stability of these compounds when they are embedded in the steel matrix. Prior APT results had shown no V-containing precipitates other than VN before or after irradiation, which motivates the use of the energy of V and N dissolved in an Fe matrix as the reference energies. Their formation and dissolution will depend on the surrounding nearby solutes rather than pure elemental materials. The reference structures for the interstitial-type elements (C and N) consisted of 128 Fe atoms and a single interstitial atom. Both the tetrahedral and octahedral site calculations were performed, and the lowest-energy sites were selected.
@@ 3 Results
[…]
Figure 1: The lattice parameters calculated using VASP for defects in rocksalt VN as a function of concentration are compared to the experimental value (unknown stoichiometry) from TEM highlighted with a horizontal line. The effect of vacancies, interstitial, and antisites are compared in a). b) shows a convex hull calculated with over 800 enumerated structures calculated with the Orb-v1 UMLIP. The red line indicates the convex hull with diatomic N and BCC V as reference states, while the blue line is the convex hull where V and N as dilute solutes in Fe are used as reference states.
The Orb machine learning potential was tested against DFT to ensure accuracy (Table 1). The Orb potential aligns well with the DFT values for stable structures and vacancies. However, the machine learning potential does not capture the high-energy unstable structures or structures containing interstitial-type defects well. This is likely due to bias from the training data, which includes vacancy-like defects but limited interstitial-type and high-energy structures.
Table 1: The root mean squared error (RMSE) values for UMLIP values against the DFT values for point defects in VN calculated over 52 structures.
[…]
Overall 0.36 0.28 0.52 0.29
V-N anti site 0.69 0.26 0.73 0
N-V anti site 0.30 0.25 0.37 0.37
V vacancies 0.14 0.07 0.09 0.03
N vacancies 0.14 0.02 0.03 0.03
V interstitials 0.40 0.39 0.48 0.42
N interstitials 0.17 0.38 0.22 0.11
The UMLIP calculated convex hull is shown in Fig 1 b). This convex hull is in good agreement with a previously reported VN convex hull[10] in that it reproduces two key features: non-stoichiometric N-vacancy-driven rocksalt structure are predicted as ground states, and these structures have highly ordered layered-vacancy structures. However, the effect of the chosen reference states plays a role in the relative stability of N vacancies when the VN is considered to be part of the wider system with the solute reference states. When the system is referenced to V and N as dilute solutes in Fe, the hull is pulled below the ground states with higher N vacancy concentration. Structures with lower N vacancy concentrations remain on the hull. These calculations combined with the observed contracted lattice parameter suggest that N vacancies may be present in the VN precipitates in significant concentrations prior to irradiation.
@@
DFT calculations of the lattice parameters as a function solute concentration are compared to the observed lattice parameters in Fig. 3 a). Substitutions of carbon on the nitrogen sublattice increase the lattice parameters while both Fe and Cr substitutions on the vanadium sublattice are predicted to decrease the lattice parameter. Although substitutions change the lattice parameter, the effect is minimal with changes being less than 0.2%. Thus, Cr and Fe substitutions may slightly alter the lattice parameter but are not a significant drivers of the lattice contraction measured here.
Figure 3: The effect of substitutional C, Fe, and Cr on lattice parameter as calculated by DFT are plotted in a) and compared to an experimental value from a single measurement at an unknown stoichiometry (the dashed horizontal line). Three ternary phase diagrams (b) V-N-C, c) V-N-Cr, d) V-N-Fe) were calculated using Orb-v1 where all points shown are on the convex hull.
To better understand the effect of these solutes in VN before irradiation, and the effects of compositional changes during irradiation, the convex hulls (Fig. 3) of three V-N-X ternary systems were calculated with DFT where X= Cr, Fe and C. In the ternary convex hull plots, the V-N-Cr system (Fig. 3 c) shows stable Cr substitutions on the V sites, which may explain the presence of Cr in the VN precipitates as found using APT. The VN precipitates show limited C content from the APT data, as shown in Fig. 2. This does not align with the ML potential enumerated convex hull shown in Fig. 3 b), where C is found on N sites. However, it is known that C content is underestimated in APT due to its tendency for high multiple detector events [25].
The presence of Fe in VN is particularly important when considering irradiation damage in Fe-based RAFM steels. Under high-energy collision cascades, Fe atoms will be ballistically implanted and point defects will accumulate at the interface between the Fe and the VN. Therefore, because the FeN rocksalt is unstable, Fe substitution in VN could destabilise and contribute to precipitate dissolution. Unfortunately, the Fe concentration in VN is difficult to determine using APT due to the field evaporation-induced artefacts. In addition, Si was found at higher concentrations in VN with APT, and this may also be an artefact of APT due to the ease of migration of Si [26]. Fe had no stable structures on the hull, as shown in Fig. 3 d). This indicates Fe substitutions or Fe interstitial accumulation may act to destabilize the VN.
Seven more ternary systems have been investigated using the Orb interatomic potential: V-N with P, S, Mn, W, Ta, Si and B Fig 4. These additional elements have been investigated because they are present in ARAFM steel [6] or, in the case of B, in boron reduced activation ferritic-martensitic steel (BRAFM) steel [27]. Each ternary plot includes the V-N binary hull already discussed on the right most edge as well as elemental reference structures at each vertex. Additional points plotted on the ternary diagram indicate additional stable structures in the ternary system. Our study did not yield any additional rocksalt-derived structures on the convex hull for the Si and P systems (Fig. 4 e) and c)). However, the remaining S, W, Ta, Mn, and B all contain stable solute-containing structures on the hull, Fig. 4 a), b), d), f) and g). As expected, the smaller radii elements of B (Fig. 4 a) and S Fig. 4 d) occupy the N sites, while the larger W and Ta occupy the V sites. Interstitial solutes were predicted to be unstable in all of the ternary systems studied.
Figure 4: Seven ternary systems are presented as ternary phase diagrams, calculated using Orb-v1, where all points are on the convex hull. a) V-N-B, b) V-N-Mn, c) V-N-P, d) V-N-S, e) V-N-Si, f) V-N-Ta, g) V-N-W.

### v2

@@ Abstract
[…]
@@
Vanadium nitride , Reduced activation ferritic martensitic steel , Universal machine learning potential , Atom probe tomography
journal: Journal of Nuclear Materials
@@
{highlights}
APT results show that Cr, along, with other solutes, are present in these precipitates.
The small experimentally measured lattice parameter for VN precipitates can be at least partially explained by a combination of N-vacancies and solutes.
NEP and DFT results both predict sub-stoichiometric ground states in the VN system with ordered vacancy layers favoured.
Switching from bulk to local reference states changed the apparent stability of vacancies in VN.
Monte Carlo and cascades confirm the presence of ordered vacancies at high temperature and their role in mitigating irradiation damage
@@ 1 Introduction
One challenge for the commercialisation of fusion energy reactors is the development of reduced activation steels with high resistance to creep and irradiation damage that are easy to manufacture [30]. To this end, advanced reduced activation ferritic-martensitic (ARAFM) steel alloys with rocksalt-structured MX phase precipitates, such as vanadium nitride (VN), have been proposed [10, 27, 17, 25]. However, Haley et al. recently observed that VN precipitates underwent complete dissolution during a 100 dpa Fe+ ion irradiation [16]. ARAFM alloys are complex systems with multiple alloying elements in addition to precipitate phases. Thus, understanding the thermodynamic factors and role of irradiation damage in precipitate stability is both challenging and necessary.
Although alloy designers have often assumed MX precipitates to have idealized rocksalt stoichiometry and crystal structures, many rocksalt nitrides exhibit a strong preference for vacancy-driven non-stoichiometry at 0 K[2, 3, 23, 32, 14, 13]. This preferred vacancy concentration may change if the MX phase is alloyed with other elements, either during precipitation or due to ballistic mixing. Additionally, during irradiation, the preferred vacancy concentration will be disturbed and other point defects will accumulate, which can further change the stability of the precipitate. A better understanding of the role of solutes and point defects on the stability and resilience of the VN precipitate under irradiation and different temperatures is needed.
We report a combined computational and experimental approach that sought to shed light on the structure of VN precipitates in ARAFM steel prior to irradiation and the types of irradiation damage that may cause dissolution. Atomic scale modelling, combined with transmission electron microscopy (TEM) and APT results, pointed to a significant concentration of vacancies and solutes in the unirradiated precipitates. Due to the high temperature operating conditions of ARAFM steels, we performed Monte Carlo and molecular dynamics at temperature using a finetuned the universal neuroevolution potential (NEP) [12, 26]. The finetuned NEP potential was then used to study the effects of irradiation damage and temperature on stability of the rocksalt V-N-X (X=Cr, Fe, C, Si, Mn, W, Ta, B, S, P) system.
@@ 2 Methods
Samples of unirradiated ARAFM steel of the same composition and microstructure, as studied under irradiation by Haley et al, were studied with TEM and APT. The lattice parameter of VN precipitates could be measured from the series of dark-field images collected by Haley et al. For each dark-field image, a corresponding diffraction pattern was captured, and the position of the objective aperture was recorded. These were compared against a simulated pattern of VN, from which the lattice parameter was deduced. APT on the unirradiated ARAFM steel was collected via both voltage (20% pulse fraction, 200 kHz pulse frequency) and laser pulsing methods (40 pJ pulse energy, 200 kHz pulse frequency) at a temperature of 50 K. This data was analysed with APSuite 6.1. The precipitate compositions were measured using proxigrams created from 2.2% concentration isoconcentration surfaces with a bin size of 0.1 nm over three precipitates.
[…]
@@
The thermodynamic stability of the VN and VNX systems was studied by analysing the formation energy crystal structures with point defects and clusters of point defects. Thirty intrinsic point defects were calculated with DFT: V(N) vacancies, interstitials, and anti-sites. An additional 18 DFT calculations were performed for Cr and Fe substitutions on the V sublattice and C substitutions on the N sublattice. The ideal rocksalt unit cell has a single unique site for each defect type. To vary the defect concentration, a single defect was inserted in rocksalt supercells ranging in size from 8 to 256 atoms. Configurations with multiple defects of different types were not considered as the primary objective was to isolate the effect of each defect. This approach necessarily limits the analysis to non-interacting defects.
To ensure consistency in the training data, the Materials Project trajectory DFT parameters [18] were used for training data generation, as the importance of maintaining consistency has been demonstrated by Mazitov et al. [28]. The training data consisted of 2797 structures, as detailed in Table 1, containing VN, VN doped with solutes, body-centered cubic Fe with solutes and ordered structures from the materials project database [18]. The data was split into a training and testing set at a 80:20 ratio. In total 11 solutes were included; Fe, B, C, Cr, Mn, P, S, Si, Ta, Ti and W. The inclusion of 11 solutes was to enable the subsequent modelling of true chemical complexity found in ARAFM steel. However, such a large element count will necessarily lower benchmark accuracy against DFT data. These structures were generated using NepTrainKit [9] and the Integrated cluster expansion toolkit (ICET). The non-ICET enumerated structures had an atomic position rattle of 0.05 Å and a random lattice scaling of 1%. All non-enumeration structures were subsampled down by a factor of five using farthest point sampling to ensure training efficiency [11]. The enumeration structures were generated using ICET [1] to enumerate all possible configurations of defects and substitutions in the V-N-X system up to 14 atoms for VN and 11 for V-N-X with X denoting one of the 11 solute elements.
Table 1: Summary of DFT training structures
System Number of structures
VN with solutes
   Structure enumeration (up to 12 atoms) 1569
   Randomly doped VN (32 atoms) 175
VN
   Structure enumeration (up to 14 atoms) 867
   Large VN structures (64 to 96 atoms) 54
   Large VN structures with vacancies (31 to 95 atoms) 22
Other
   Solute reference states in Fe (128 to 129 atoms) 13
   Ordered phases (up to 126 atoms) 97
Total 2797
The NEP potential was fine-tuned from the NEP89 foundational model [26]. The potential was fine-tuned with the following training parameters
$\lambda_{1}=\lambda_{2}=0.2,$ (1)
$\lambda_{E}=10,$ (2)
$\lambda_{F}=\lambda_{V}=1,$ (3)
where $\lambda_{1}$ and $\lambda_{2}$ are the $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ norm regularisations respectively. $\lambda_{E}$, $\lambda_{F}$ and $\lambda_{V}$ are the energy, forces and virials training weights respectively. In addition, the training was run for 5000 generations with a population size of 20 and a batch size of 2244, chosen to match the entire training dataset size. The model performance is presented in Table 2. The root mean squared error (RMSE) values improved during finetuning and reached comparable errors to large message passing neural networks, such as orb-v3 [33], but with a substantially lower inference computational cost of approximately three orders of magnitude. The testing dataset parity plots for energy, forces, stresses and virials are included in the supplementary information in fig. 9. This reduced computational cost was prioritised over only minimal errors to be capable of performing molecular dynamics of large systems. Crucially the NEP potential after finetuning was able capture the same vacancy ordered structures on the convex hull as DFT indicating its suitability.
Table 2: Model performance metrics
Metric Error
RMSE test energy 0.068 eV
RMSE test forces 0.577 eV/Å
RMSE test virials 0.263 eV/atom
RMSE VN hull distance 0.242 eV/atom
RMSE VN hull distance 0.093 eV/atom
   (within 0.2 eV)
[…]
@@
Structures were calculated with the NEP as follows. First, an ideal conventional unit cell was initialized with ASE. Supercells were then enumerated using ICET [1]. The structure enumeration was broken into different series for each defect and substitution type. All possible unique structures containing the defect type were enumerated using ICET to a maximum of 14 atoms for the rocksalt structure. The NEP potential was used to calculate formation enthalpies for defect in VN. The defect types consisted of both V-N anti-sites, both V and N interstitials and both V-N vacancies. While this was not an exhaustive search of all phases that could be included in the system, it does include the unique single type defect in the VN rocksalt system and each element dissolved in BCC Fe, i.e. the two phases observed in experimental analysis.
@@
While formation energies are typically calculated relative to the elements standard crystal structures, we have chosen to use the energy of elements as solutes in (BCC) iron. This was chosen because we are most interested in the stability of these compounds when they are embedded in the steel matrix. Prior APT results had shown no V-containing precipitates other than VN before or after irradiation, which motivates the use of the energy of V and N dissolved in an Fe matrix as the reference energies. Their formation and dissolution will depend on the surrounding nearby solutes rather than pure elemental materials. The reference structures for the interstitial-type elements (C and N) consisted of 128 Fe atoms and a single interstitial atom. Both the tetrahedral and octahedral site calculations were performed, and the lowest-energy sites were selected.
To investigate the ordering of vacancies at the high temperatures targeted by ARAFM steel of around 900 K an on-lattice Monte Carlo approach was taken to address the 0 K limitation of DFT. A cubic supercells containing 2744 atoms were used with different initial compositions of VN with solutes and vacancies. The vacancies are represented as a unique species labelled X. A Monte Carlo algorithm was then used to find a representative low energy configuration for each of the given conditions. The simulations were carried out on-lattice to enable the swapping of vacancies, which remains difficult for off-lattice simulations.
The Monte Carlo algorithm used is as follows. Where $E_{i}$ is the energy before a swap, $E_{f}$ is the energy after a swap, $k_{B}$ is the Boltzmann constant and $T$ is temperature.
$\Delta E=E_{f}-E_{i}$ (5)
[…]
To carry out the Monte Carlo simulations the software package Clease [8] was used with the finetuned NEP potential. The simulations were run for a total of 70,000 attempted swaps for 3 conditions; VN at 933 K with different initial vacancy concentrations, VN with 25% vacancies on the N sublattice at temperatures between 0 and 2000 K and VN with solutes and vacancies representative of the APT concentration of 5% vacancy and 5% C substitutions on the N sublattice and 12% Cr, 1% Mn, 1% W, 2% Si and 5% Fe on the V sublattice.
To quantify the vacancy ordering, the short range order (SRO) parameter was quantified for each neighbour shell, k, from 1 to 10. The SRO was calculated by
$\alpha_{ij}^{(k)}=1-\frac{n_{ij}^{(k)}}{2c_{j}}$ (7)
where $\alpha_{ij}^{(k)}$ is the SRO parameter between elements $i$ and $j$ for neighbour shell $k$, $n_{ij}^{(k)}$ is the number of atoms of type $j$ neighbouring atoms of type $i$ in shell $k$, $c_{j}$ is the concentration of atoms of type $j$ in the bulk, and the factor of 2 scales the SRO to one sublattice.
The final structures from the following Monte Carlo runs at 933 K; (5%, 14%, 26%, 50%) vacancy concentration and the APT representative concentration were used as the starting configurations for primary knock-on atom (PKA) cascades. The PKA cascades were carried out three times for each of the [100], [110] and [111] directions and for each starting structure. The PKA energy was set to 1200 eV and the temperature of the simulation was set to 900 K. The structure were first relaxed using the FIRE algorithm for 2000 steps and then thermalised using the stochastic rescaling NPT ensemble [4] for between 5 ps to 14 ps. The thermalisation time was varied between each run of the same structure to result in a different random seed for the cascade simulation. After thermalisation, the central N atom was selected and given a kinetic energy of 1200 eV. The NVE ensemble was used in the central region and a Nose-Hoover chain (NHC) thermostat was set to a 10 Å thick boundary at the outer edge of the supercell. The NHC thermostat was given a target temperature of 900 K with a temperature coupling constant of 200 K. The timestep was set at 1 fs with a maximum displacement per step of 0.015 Å.
To quantify the damage following the cascades the change in stored energy was calculated. This was used instead of the defect count that would be unreliable in this case due to the presence of vacancy defects in the initial structures. The stored energy, $\Delta E_{\mathrm{stored}}$, is given by the relaxed energy after the cascade, $E_{\mathrm{final}}$ minus the relaxed energy before the cascade $E_{\mathrm{initial}}$, or simply
$\Delta E_{\mathrm{stored}}=E_{\mathrm{final}}-E_{\mathrm{initial}}.$ (8)
@@ 3 Results
The experimental lattice parameter for the VN precipitates as measured with TEM was $4.035\r${A}, a smaller measurement than the previously reported ($4.1287\r${A}) [24] and much smaller than the calculated lattice parameter for the pristine VN rocksalt structure. This measurement is compared to the lattice parameter of VN as a function of the concentration of intrinsic point defects (vacancies, antisites, and interstitials) as calculated with DFT in Figure 1a). Vacancies on both the vanadium or nitrogen sublattices both lead to a reduction in the lattice parameter (Figure 1a) of up to 2%, whereas the anti-sites and interstitials increase the lattice parameter by 4% and 8% respectively.
Figure 1: The lattice parameters calculated using VASP for defects in rocksalt VN as a function of concentration are compared to the experimental value (unknown stoichiometry) from TEM highlighted with a horizontal line. The effect of vacancies, interstitial, and antisites are compared in a). The effect of substitutional C, Fe, and Cr on lattice parameter as calculated by DFT are plotted in b) and compared to an experimental value from a single measurement at an unknown stoichiometry (the dashed horizontal line).
The NEP calculated convex hull is shown in Fig 2. This convex hull is in good agreement with a previously reported VN convex hull[32] in that it reproduces two key features: non-stoichiometric N-vacancy-driven rocksalt structures are predicted as ground states, and these structures have highly ordered layered-vacancy structures. However, the effect of the chosen reference states plays a role in the relative stability of N vacancies when the VN is considered to be part of the wider system with the solute reference states. When the system is referenced to V and N as dilute solutes in Fe, the hull is pulled below the ground states with higher N vacancy concentration. Structures with lower N vacancy concentrations remain on the hull. These calculations combined with the observed contracted lattice parameter suggest that N vacancies may be present in the VN precipitates in significant concentrations prior to irradiation.
Figure 2: The convex hull calculated with 6519 enumerated structures calculated with the finetuned NEP potential. The red line indicates the convex hull with molecular N and BCC V as reference states, while the blue line is the convex hull where V and N as dilute solutes in Fe are used as reference states.
DFT calculations of the lattice parameters as a function solute concentration are compared to the observed lattice parameters in Fig. 1 b). Substitutions of carbon on the nitrogen sublattice increase the lattice parameters while both Fe and Cr substitutions on the vanadium sublattice are predicted to decrease the lattice parameter. Although substitutions change the lattice parameter, the effect is minimal with changes being less than 0.2%. Thus, Cr and Fe substitutions may slightly alter the lattice parameter but neither are a significant driver of the lattice contraction measured here.
@@
Ordered vacancy layers have been confirmed in this work at zero temperature and with respect to solute reference state, as shown in the VN convex hull in Fig. 2. However, to investigate the effect of temperature on-lattice Monte Carlo simulations were conducted. The SRO between vacancy-vacancy at 933 K in the first to tenth nearest neighbour shells (the sites are illustrated in Fig. 4 c)) as a function of vacancy concentration was found to trend to zero and random ordering with increasing vacancy concentrations, as shown in Fig. 4 a). In addition, a high SRO parameter can be seen for the shell numbers 2 and 4 highlighting less than random occupations. The sixth and seventh neighbour shells are found to have the lowest SRO parameters, although the ordering decreases with increasing vacancy concentration. The effect of temperature was subsequently investigated with a fixed vacancy concentration of 25%. As with the vacancy concentration, increasing temperature was found to decrease ordering, as shown in Fig. 4 b). It is noted that significant ordering still exists at 933 K for shells 2 and 6. Shell 4 exhibits unusual behaviour likely a result of fluctuations between many configurations close in energy. The APT representative composition was also investigated at 933 K and the SRO as function of shell number is shown in Fig. 4 d). The Cr, Mn, and Si show the most ordering with all solutes having weaker ordering at larger distances.
Figure 4: The ordering present taken from the final states from on-lattice Monte Carlo simulations. a) shows the SRO parameters as a function of initial vacancies (X) on the N sublattice at a temperature of 933 K. b) shows the SRO parameters as a function of temperature at a vacancy concentration of 25%. c) an illustration of nearest neighbour shells of N atoms in a slice in the [111] plane viewed down the [111] direction. d) shows the SRO parameters as a nearest neighbour shells including solutes at a temperature of 933 K.
The vacancy ordering behaviour showed stark differences with VN and VN with the added solutes. The vacancies on the N sublattice has been shown to order along the [111] plane. However, with the introduction of solute atoms, most notably large atoms such as W, this ordering is broken down and the formation of clustering like behaviour was observed, as shown in Fig. 5. In particular, in Fig. 5 a) and b) the vacancies form into vacancy-W-Si-Mn clusters. This is likely partially because the presence of vacancies mitigating the strain induced from the large W atoms.
Figure 5: The on-lattice Monte Carlo atomic configurations taken from the final state of a 5% vacancy concentration at 933 K with and without solutes. a) and c) show the vacancies in $\mathrm{VN}_{0.95}$ condition coloured by species, a), or by local vacancy concentration (within 3 Å). b) and d) show the vacancies in $\mathrm{VN}_{0.95}$ containing solutes condition coloured by species, a), or by local vacancy concentration (within 3 Å).
To understand the dynamics of irradiation damage, PKA collision cascades were performed on five conditions; VN, VN with a (5%, 14%, 26%, 50%) pre-existing vacancy concentration and the APT representative concentration. All starting structures were taken from the final state of the Monte Carlo simulations. As a result structures containing vacancies started with ordered vacancies, and the APT representative structure contained vacancy-solute clusters. The lower initial vacancy concentration conditions of VN0.86, VN0.95 and VN0.74 had a lower increase in stored energy when compared with the pristine VN, as shown in Fig. 6 a). The higher vacancy concentration VN0.5 was found to have a larger increase in stored energy over the pristine case. Finally, the APT composition had an apparent protective effect due to the presence of 5% vacancies and an apparent detrimental effect due to the addition of solutes. The two competing effects result in a comparable stored energy to the pristine condition. The Frenkel pair count over time, given by Fig. 6 b), shows a clear relationship between vacancy number and final number of Frenkel pairs. However, this is caused by the measurement including the movement of defect as opposed the their creation. Therefore, the stored energy is a more applicable metric that include the increase in energy from both the disordering of defects as well as the creation of new defects.
Figure 6: The damage introduced from collision cascades. a) showing the change in stored energy for each condition. b) The Frenkel pair count over time for each condition calculated from the initial frame starting condition.